# Intervention Efficiency and Perturbation Validation Framework: Capacity-Aware and Robust Clinical Model Selection under the Rashomon Effect

### v1

@@ Abstract
In clinical machine learning, the coexistence of multiple models with comparable performance—a manifestation of the Rashomon Effect—poses fundamental challenges for trustworthy deployment and evaluation. Small, imbalanced, and noisy datasets, coupled with high-dimensional and weakly identified clinical features, amplify this multiplicity and make conventional validation schemes unreliable. As a result, selecting among equally-performing models becomes uncertain, particularly when resource constraints and operational priorities are not considered by conventional metrics like F1 score. To address these issues, we propose two complementary tools for robust model assessment and selection: Intervention Efficiency (IE) and the Perturbation Validation Framework (PVF). IE is a capacity-aware metric that quantifies how efficiently a model identifies actionable true positives when only limited interventions are feasible, thereby linking predictive performance with clinical utility. PVF introduces a structured approach to assess the stability of models under data perturbations, identifying models whose performance remains most invariant across noisy or shifted validation sets. Empirical results on synthetic and real-world healthcare datasets show that using these tools facilitates the selection of models that generalize more robustly and align with capacity constraints, offering a new direction of tackling the Rashomon Effect in clinical settings.
@@ 1 Introduction
High dimensionality further increases overfitting and obscures relevant predictors (berisha2021digital; bommert2022benchmark), motivating feature selection or dimensionality reduction (jolliffe2016principal; jia2022feature). Still, models like logistic regression, SVMs, k-nearest neighbors, and random forests can achieve similar accuracies while relying on distinct feature sets (lukaszuk2024stability). This phenomenon, called the Rashomon Effect (breiman2001statistical; rudin2024amazing; xin2022exploring), reveals that many “equally good” but mutually-exclusive models may coexist, complicating interpretation and clinical trust.
@@
To address these challenges, we propose two complementary tools: Intervention Efficiency (IE) and the Perturbation Validation Framework (PVF). IE quantifies how effectively a model identifies minority cases when intervention capacity is limited, making the precision–recall trade-off explicit and clinically meaningful. PVF stress-tests models across perturbed validation sets and selects the one with the most stable performance, emphasizing robustness and interpretability. While introduced alongside IE, PVF is fully compatible with conventional metrics such as F1 score or accuracy. Experiments on synthetic and real medical datasets show that PVF guides model selection toward greater generalization under both IE and traditional metrics. Together, IE and PVF provide a resource-aware, robustness-focused evaluation framework that facilitates model selection when the Rashomon Effect appears in small, imbalanced healthcare datasets (abdullah2021review; Christodoulou2019; rudin2019stop; stiglic2020interpretability).
@@ 2 Related Work
Small, imbalanced datasets create high variance, overfitting, and instability in model and feature selection (cawley2010overfitting; zhang2023model; xin2022exploring), making it difficult to select one reliable model among many, especially when the Rashomon Effect appears. In low-sample settings, small data perturbations can cause drastic changes in model structure and predictions, especially for unstable learners such as decision trees or stepwise regression (breiman1996bagging; breiman1996heuristics; morozova2015comparison; meinshausen2010stability; dietterich2000ensemble). Under class imbalance, models may favor majority classes or overfit rare ones (chawla2002smote; he2009learning; buda2018systematic; vandenGoorbergh2022harm).
@@ 3.1 Intervention Efficiency (IE)
In the following proposition, we show that the IE of a model $f$ depends only on its precision $p$, recall $r$, the population prevalence $\pi$, and the intervention capacity $\gamma$.
@@
Assume that there is a model $f$ that can predict binary labels and it is used for model-guided intervention. The problem has an overall positive rate $\pi$ and only a fraction $\gamma$ of the whole population can be intervened due to limited capacity. The Intervention Efficiency of $f$ on this problem under intervention capacity $\gamma$, denoted as $\mathrm{IE}_{\gamma}(f)$, satisfies
@@
We refer readers to Appendix A for the the detailed proof of Proposition 3.1. In practice, we can approximate the true $\pi$ by the positive rate $\hat{\pi}$ in the specific set of samples to which the model is applied. In other words, for instance, when estimating IE on a validation set, even with the whole dataset that can provide a ”more accurate estimation” for $\pi$, we should use the $\hat{\pi}$ from the specific validation set rather than that from the whole dataset. The reason is that IE is tied to precision and recall, two values depend not only on the model but also largely on the set of samples that yields them, especially in small sample cases. In contrast, in large sample cases, $\hat{\pi}$ has been a good estimate of true $\pi$ given enough samples. In sum, it is both reasonable and essential to estimate $\pi$, $p$ and $r$ on the same set to ensure they are aligned, since they are jointly dependent on each other. Practically, when computing IE on a specific set of samples $\mathcal{D}$, it not only depends on a model $f$ and a $\gamma$ of interest, but also on $\mathcal{D}$, where $\hat{\pi}$, $\hat{p}$ and $\hat{r}$ (estimators of $\pi$, $p$ and $r$) are derived. Consequently, we can rewrite the formula for IE in real practice as:
@@ 3.2 Perturbation Validation Framework (PVF)
Finally, select $f^{\star}\in\arg\max_{f\in\mathcal{F}}A_{f}$, as shown in Figure 5. The complete implementation pseudocode of combining PVF with any evaluation metric $\mathrm{Eval}$, denoted as PVF($\mathrm{Eval}$), appears in Algorithm 1.
@@
    $\mathcal{A}:\mathbb{R}^{M}\to\mathbb{R}$: aggregation function (e.g., a lower quartile).
Since no retraining is required, the per‐model computational complexity is $\mathcal{O}(Mknd)$, and across all $Q$ models it is $\mathcal{O}(QMknd)$.
Algorithm 1 PVF(Eval) Model Selection
@@ 4 Experiments
This section outlines the objectives, general setups, and overall findings of our experiments on both synthetic and real datasets. Detailed descriptions, full quantitative results, and sensitivity analyses are provided in the Appendix. Specifically, Appendix C present the detailed setups and results for the synthetic dataset experiments, and Section 5.2 presents the detailed setups for real datasets, respectively, while Appendix E reports a comprehensive hyperparameter sensitivity analysis.
4.1 Experimental Research Questions
@@
    On real clinical datasets, does PVF choose better models than the traditional approach under the same metrics used in the synthetic dataset experiments (Section 5.2)?
@@ 4.2 General Experimental Setup
After selection, the chosen model is either assessed based on ground truth information or when it is not available, as an approximation of it, on a large external test set using the same metric as in validation. To improve reliability and reduce variance from any single split, each configuration is repeated extensively. Model classes and other dataset-specific design choices are kept simple and interpretable to mimic clinical use and are described in subsequent subsections.
5 Overview of Experimental Results
@@
5.1 Synthetic Dataset Experiments
As shown in Figure 2, PVF consistently outperforms the traditional validation method across most configurations (the detailed configurations are elaborated in Section C.1). When $\gamma=0.1$ and $\gamma=0.3$, PVF achieves superior selections in approximately 90% of all configurations, indicating PVF’s advantage under limited intervention capacities. Although the advantage narrows as $\gamma$ increases ($\gamma\geq 0.7$), PVF still maintains a clear lead.
Figure 2: Synthetic experiments: percentage of configurations where PVF outperforms the traditional method across all metrics. Positive Difference shows the percentage of configurations PVF more often selects the best model than the traditional method does across all repetitions.
5.2 Real-World Dataset Experiments
In this subsection, similar to the synthetic dataset experiment, we also present the experimental results in the form of a comparison between the performance of PVF under $\sigma$ (defined as in Section 3.2) tuning and that of the traditional method, respectively for the cervical cancer (wolberg1993breast) and breast cancer (fernandes2017cervical) datasets. We refer the readers to Appendix D for the detailed setups of this part of experiments.
@@ 5.2 Real-World Dataset Experiments
For the cervical cancer dataset, PVF achieves its best results with a small perturbation for IE ($\sigma=0.01$) and a near-zero perturbation for F1 Score ($\sigma=10^{-6}$), as shown in Figure 3. At the most stringent intervention capacity ($\gamma=0.1$), PVF selects an externally better model in 60.0% of times, compared to 26.7% for the traditional method and 13.3% ties. As $\gamma$ increases to 0.3–0.9, PVF continues to win more frequently, though the margin narrows: 43.3–46.7% times PVF wins versus 33.3–36.7% the traditional method does, with around 20% ties across these settings. For F1, PVF still leads with 50.0% wins, compared to 33.3% the traditional method does and 16.7% ties. These results indicate that once $\sigma$ is tuned, PVF consistently yields more reliable model selection than the traditional method, with its strongest advantage at $\gamma=0.1$, where false positives are extremely costly.
Figure 3: Best Performance for Cervical Cancer Dataset under Different Evaluation Metrics
@@
Figure 4: Best Performance for Breast Cancer Dataset under Different Evaluation Metrics
@@
5.3 Brief Conclusion of Experiments
@@ 6 Conclusion
This work addresses a persistent challenge in clinical machine learning: selecting a single, interpretable model when data are limited, class distributions are imbalanced, and measurements are noisy, potentially aggravating the Rashomon Effect. We propose Intervention Efficiency (IE), a capacity-aware evaluation metric that quantifies the expected benefit of a model when only a fixed fraction of individuals can receive an intervention, and Perturbation Validation Framework (PVF) that prioritizes candidate models whose performance remains strong under perturbations of the validation set. Together, IE and PVF facilitate model selection with operational constraints and robustness requirements that are typical of clinical deployment.
Empirical studies spanning an exhaustive synthetic enumeration and two clinical datasets indicate that the PVF more oftern identifies models that generalize better out of sample than a conventional single-split selection. The gains are most pronounced in the regimes with a low intervention capacity and when F1 score is the metric of interest.
As for limitations and future work, the selection and calibration of perturbation noise for PVF are consequential and may require domain-expert input to ensure that perturbations faithfully reflect clinical measurement variability and data quality. Otherwise, PVF may fail to outperform the traditional validation approach. Therefore, developing principled elicitation methods (e.g., protocol- or instrument-informed priors) and practical defaults for common data types would be crucial to improve applicability. On the theory side, there is substantial scope for formal analysis of PVF, including consistency guarantees and sensitivity to different performance metrics, especially the newly proposed IE. Such results clarify applicable scenarios defined by data scale, noise, and resource limits. For IE, extensions to multiclass outcomes, integration of explicit cost, and consideration of fairness constraints remain important directions. Overall, coupling PVF with IE offers a practical, conceptual pathway and a future direction for navigating the Rashomon Effect in clinical machine learning, helping translate model multiplicity from a source of uncertainty into a framework for trustworthy and resource-conscious model selection.
@@
We present the derivation in five steps including the analysis of the model’s capability, the two operating regimes, the expression unification for the two regimes, and the transition from counting-based expression to the final ratio-based one. For clarity, we denote $\alpha$ as the total number of positives in the population, $\beta$ as the population size, and $c$ as the number of individuals that can be intervened. Specifically,
@@
    where $\beta_{\mathrm{rem}}$ denotes the size of the residual population after removing the $\tfrac{\alpha r}{p}$ intervened, model-flagged individuals; $\alpha_{\mathrm{rem}}$ is the number of true positives remaining in that residual population; and $\pi_{\mathrm{rem}}$ is the corresponding positive rate within this residual pool. We assume that the predictive model is uninformative on this residual pool. Therefore, in stage 2, the remaining $c_{\mathrm{rem}}$ interventions are allocated uniformly at random, yielding

### v6

@@ Abstract
In clinical machine learning, the coexistence of multiple models with comparable performance—a manifestation of the Rashomon Effect—poses fundamental challenges for trustworthy deployment and evaluation. Small, imbalanced, and noisy datasets, coupled with high-dimensional and weakly identified clinical features, amplify this multiplicity and make conventional validation schemes unreliable. As a result, selecting among equally-performing models becomes uncertain, particularly when resource constraints and operational priorities are not considered by conventional metrics like F1 score. To address these issues, we propose two complementary tools for robust model assessment and selection: the Intervention Efficiency (IE) and the Perturbation Validation Framework (PVF). IE is a capacity-aware metric that quantifies how efficiently a model identifies actionable true positives when only limited interventions are feasible, thereby linking predictive performance with clinical utility. PVF introduces a structured approach to assess the stability of models under data perturbations, identifying models whose performance remains most invariant across noisy or shifted validation sets. Empirical results on synthetic and real-world healthcare datasets show that using these tools facilitates the selection of models that generalize more robustly and align with capacity constraints, offering a new direction of tackling the Rashomon Effect in clinical settings.
Codehttps://github.com/YuwenZhang-Peter/PVF-IE
@@ 1 Introduction
High dimensionality further increases overfitting and obscures relevant predictors (berisha2021digital; bommert2022benchmark), motivating feature selection or dimensionality reduction (jolliffe2016principal; jia2022feature). Still, models like logistic regression, SVMs, k-nearest neighbors, and random forests can achieve similar accuracies while relying on distinct feature sets (lukaszuk2024stability). This phenomenon, called the Rashomon Effect (breiman2001statistical; rudin2024amazing; xin2022exploring), reveals that many “equally good” but mutually-exclusive models may coexist. This multiplicity usually complicated interpretation and clinical trust.
@@
To address these challenges, we propose two complementary tools: the Intervention Efficiency (IE) and the Perturbation Validation Framework (PVF). IE quantifies how effectively a model identifies minority cases when intervention capacity is limited, making the precision–recall trade-off explicit and clinically meaningful. PVF stress-tests models across perturbed validation sets and selects the one with the most stable performance, emphasizing robustness and interpretability. While introduced alongside IE, PVF is fully compatible with conventional metrics such as F1 score or accuracy. Experiments on synthetic and real medical datasets show that PVF guides model selection toward greater generalization under both IE and traditional metrics. Together, IE and PVF provide a resource-aware, robustness-focused evaluation framework that facilitates model selection when the Rashomon Effect appears in small, imbalanced healthcare datasets.
@@ 2 Related Work
Small, imbalanced datasets create high variance, overfitting, and instability in model and feature selection (cawley2010overfitting; zhang2023model; xin2022exploring), making it difficult to select one reliable model among many, particularly when the Rashomon Effect appears and causes multiplicity in models. In low-sample settings, small data perturbations can cause drastic changes in model structure and predictions, especially for unstable learners such as decision trees or stepwise regression (breiman1996bagging; breiman1996heuristics; morozova2015comparison; meinshausen2010stability; dietterich2000ensemble). Under class imbalance, models may favor majority classes or overfit rare ones (chawla2002smote; he2009learning; buda2018systematic; vandenGoorbergh2022harm).
@@ 3.1 Intervention Efficiency (IE)
In the following proposition, we show that the IE of a model $f$ depends only on its precision $p$, recall $r$, the prevalence rate $\pi$ (i.e., the percentage of positive cases in the population), and the intervention capacity $\gamma$.
@@
Assume that there is a model $f$ that can predict binary labels and it is used for model-guided intervention. The problem has an overall prevalence rate $\pi$ and only a fraction $\gamma$ of the whole population can be intervened due to limited capacity. The Intervention Efficiency of $f$ on this problem under intervention capacity $\gamma$, denoted as $\mathrm{IE}_{\gamma}(f)$, satisfies
@@
We refer readers to Appendix A for the detailed proof of Proposition 3.1. In practice, we can approximate the true $\pi$ by the prevalence rate $\hat{\pi}$ in the specific set of samples to which the model is applied. For instance, when estimating IE on a validation set, even with the whole dataset that can provide a ”more accurate estimation” for $\pi$, we should use the $\hat{\pi}$ from the specific validation set rather than the prevalence rate from the whole dataset. The reason is that IE is tied to precision and recall, two values depending not only on the model but also largely on the set of samples that yields them, especially in small sample cases. In contrast, in large sample cases, $\hat{\pi}$ has been a good estimate of true $\pi$ given enough samples. In sum, it is both reasonable and essential to estimate $\pi$, $p$ and $r$ on the same set to ensure they are aligned, since they are jointly dependent on each other. Practically, when computing IE on a specific set of samples $\mathcal{D}$, it not only depends on a model $f$ and a $\gamma$ of interest, but also on $\mathcal{D}$, where $\hat{\pi}$, $\hat{p}$ and $\hat{r}$ (estimators of $\pi$, $p$ and $r$) are derived. Consequently, we can rewrite the formula for IE in real practice as:
@@ 3.2 Perturbation Validation Framework (PVF)
Finally, select $f^{\star}\in\arg\max_{f\in\mathcal{F}}A_{f}$, as shown in Figure 2. The complete implementation pseudocode of combining PVF with any evaluation metric $\mathrm{Eval}$, denoted as PVF-$\mathrm{Eval}$, appears in Algorithm 1.
@@
    $\mathcal{A}:\mathbb{R}^{M}\to\mathbb{R}$: aggregation function (e.g., a lower quantile).
Since no retraining is required, the per‐model computational complexity is $\mathcal{O}(Mknd)$, and across all $Q$ models it is $\mathcal{O}(QMknd)$. We refer readers to Appendix B for theoretical guarantees of PVF.
Algorithm 1 PVF-$\mathrm{Eval}$ Model Selection
@@ 4 Experiments
This section outlines the objectives, general setups, and overall findings of our experiments on both synthetic and real datasets. Detailed descriptions, full quantitative results, and sensitivity analyses are provided in the Appendix. Specifically, Appendix C present the detailed setups and results for the synthetic dataset experiments; Appendix D presents the detailed setups for real datasets; Appendix E reports a comprehensive hyperparameter sensitivity analysis.
4.1 Research Questions
@@
    On real clinical datasets, does PVF select models that perform better than those selected by the traditional approach under the same metrics used in the synthetic dataset experiments (Section 4.3)?
@@ 4.2 General Experimental Setup
After selection, the chosen model is either assessed based on ground truth information or when it is not available, as an approximation of it, on a large external test set using the same metric as in validation. To improve reliability and reduce variance from any single split, each configuration is repeated extensively. Model classes and other dataset-specific design choices are kept simple and interpretable to mimic clinical use (stiglic2020interpretability), and are described in subsequent subsections.
4.3 Overview of Experimental Results
@@
Synthetic Dataset Experiments
As shown in Figure 3, PVF consistently outperforms the traditional validation method across most configurations (the configurations are elaborated in Section C.1). When $\gamma=0.1$ and $\gamma=0.3$, PVF achieves superior selections in approximately 90% of all configurations, indicating PVF’s advantage under limited intervention capacities. Although the advantage narrows as $\gamma$ increases ($\gamma\geq 0.7$), PVF still maintains a clear lead.
Figure 3: Percentage of configurations in synthetic experiments where PVF outperforms the traditional method across all metrics. ”% Positive Difference” shows the percentage of configurations in which PVF more often selects the best model than the traditional method does across all repetitions.
Real-World Dataset Experiments
In this subsection, similar to the synthetic dataset experiment, we also present the experimental results in the form of a comparison between the performance of PVF under $\sigma$ (defined as in Section 3.2) tuning and that of the traditional method, respectively for the cervical cancer (wolberg1993breast) and breast cancer (fernandes2017cervical) datasets. We refer the readers to Appendix D for the detailed setups of this part of experiments, including exactly how we calculate the performance gaps between PVF and the traditional method with repetitive experiments.
@@
For the cervical cancer dataset, PVF achieves its largest performance gap with the traditional method with a small perturbation for IE ($\sigma=0.01$) and a near-zero perturbation for F1 Score ($\sigma=10^{-6}$), as shown in Figure 4. At the most stringent intervention capacity ($\gamma=0.1$), PVF selects an externally better model in 60.0% of times, compared to 26.7% for the traditional method and 13.3% ties. As $\gamma$ increases to 0.3–0.9, PVF continues to win more frequently, though the margin narrows: 43.3–46.7% times PVF wins versus 33.3–36.7% the traditional method does, with around 20% ties across these settings. For F1, PVF still leads with 50.0% wins, compared to 33.3% the traditional method does and 16.7% ties. These results indicate that once $\sigma$ is tuned, PVF consistently yields more reliable model selection than the traditional method, with its strongest advantage at $\gamma=0.1$, where false positives are extremely costly.
@@
Brief Conclusion of Experiments
@@
Figure 4: Largest performance gaps of PVF on the cervical cancer dataset under different evaluation metrics, when $\sigma$ is tuned. PVF consistently selects better external models than the traditional method across all $\gamma$ values, with its strongest advantage at $\gamma=0.1$.
Figure 5: Largest performance gaps of PVF on the breast cancer dataset under different evaluation metrics, when $\sigma$ is tuned. PVF tends to outperform the traditional method across lower $\gamma$ values, specifically when $\gamma\in\{0.1,0.3\}$, while tying with the traditional method at higher $\gamma$ values.
@@ 5 Conclusion
This work aim to address a persistent challenge in clinical machine learning: selecting a single, interpretable model when data are limited, class distributions are imbalanced, and measurements are noisy, potentially aggravating the Rashomon Effect. We propose Intervention Efficiency (IE), a capacity-aware evaluation metric that quantifies the expected benefit of a model when only a fixed fraction of individuals can receive an intervention, and Perturbation Validation Framework (PVF) that prioritizes candidate models whose performance remains strong under perturbations of the validation set. Together, IE and PVF facilitate model selection under operational constraints and robustness requirements that are typical of clinical deployment.
Empirical studies spanning an exhaustive synthetic enumeration and two clinical datasets indicate that the PVF more often identifies models that generalize better out of sample than a conventional single-split selection does. The gains are most pronounced in the regimes with a low intervention capacity and when F1 score is the metric of interest.
As for limitations and future work, the selection and calibration of perturbation noise for PVF are consequential and may require domain-expert input to ensure that perturbations faithfully reflect clinical measurement variability and data quality. Otherwise, PVF may fail to outperform the traditional validation approach. Therefore, developing principled elicitation methods (e.g., protocol- or instrument-informed priors) and practical defaults for common data types would be crucial to improve applicability. On the theory side, there is substantial scope for formal analysis of PVF, including consistency guarantees and sensitivity to different performance metrics, especially the newly proposed IE. Such results would clarify applicable scenarios defined by data scale, noise, and resource limits. For IE, extensions to multiclass outcomes, integration of explicit cost, and consideration of fairness constraints remain important directions. Overall, coupling PVF with IE offers a practical, conceptual pathway and a future direction for navigating the Rashomon Effect in clinical machine learning, helping translate model multiplicity from a source of uncertainty into a framework for robust and resource-conscious model selection.
6 Acknowledgments
This work was supported by the Data+X (2024) project at DKU. We would also like to thank Prof. Dongmian Zou for valuable discussions on the theoretical aspects of the Intervention Efficiency and the Perturbation Validation Framework, and Prof. Chenkai Wu for his insightful suggestions on the Intervention Efficiency that helped us connect the proposed metric to practical interpretations.
@@ 6 Acknowledgments
We present the derivation in five steps including the analysis of the model’s capability, the two operating regimes, the expression unification for the two regimes, and the transition from counting-based expression to the final ratio-based one. All notations and terms follow the same definitions in Proposition 3.1. For clarity, we additionally denote $\alpha$ as the total number of positives in the population, $\beta$ as the population size, and $c$ as the number of individuals that can be intervened. Specifically,
@@
    where $\beta_{\mathrm{rem}}$ denotes the size of the residual population after removing the $\tfrac{\alpha r}{p}$ intervened, model-flagged individuals; $\alpha_{\mathrm{rem}}$ is the number of true positives remaining in that residual population; and $\pi_{\mathrm{rem}}$ is the corresponding prevalence rate within this residual pool. We assume that the predictive model is uninformative on this residual pool. Therefore, in stage 2, the remaining $c_{\mathrm{rem}}$ interventions are allocated uniformly at random, yielding